# EpochCore: Digital Hardware Accelerator For Structured State-Space Models

### v1

@@ Abstract
In this paper we introduce a specialized hardware accelerator,  EpochCore [^1] Epoch does not refer to training iterations, instead it is an acronym., for accelerating SSMs. EpochCore is based on systolic arrays (SAs) and is designed to enhance the energy efficiency and throughput of inference of SSM-based models for long-range sequence tasks. Within the SA, we propose a versatile processing element (PE) called LIMA-PE to perform traditional and specialized MAC operations to support traditional DNNs and SSMs. To complement the EpochCore microarchitecture, we propose a novel dataflow, ProDF, which enables highly efficient execution of SSM-based models. By leveraging the LIMA-PE microarchitecture and ProDF, EpochCore achieves on average 250×~250$\times$250 × gains in performance and 45×~45$\times$45 × improvement in energy efficiency, at the expense of 2×2$\times$2 × increase in area cost over traditional SA-based accelerators, and around 2,000$\sim$ 2,000$\times$ 2 , 000 × improvement in latency/inference on LRA datasets compared to GPU kernel operations.
@@ I Introduction
To address this limitation, recent research has explored alternative architectures based on State Space Models (SSMs) [25], which offer a promising framework for scalable sequence modeling. The Linear State Space Layer (LSSL) [10] demonstrated early improvements over both RNNs and Transformers on simpler sequence tasks. Further enhancements to SSMs introduced structured state transition matrices—e.g., in S4 [9], S5 [27], DSS [11], and H3 [6]—to boost both accuracy and efficiency. More recent models, such as Liquid-S4 [13, 12], GSS [23], and Mamba [7, 21], integrate time-varying parameters and gating mechanisms, achieving state-of-the-art results on long-range benchmarks such as the Long Range Arena (LRA) [29]. These benchmarks encompass a wide range of applications, including 500+ token language tasks, 1000+ time steps for time-series data, and hundreds of video frames. For example, Gu et al. [9] report that S4 achieves a 5.19$\times$ speedup and 0.091$\times$ memory usage compared to Transformers on 4K-length inputs, while also delivering over 60$\times$ throughput gains in token processing. SSM is not suited for non-sequential data such as grid-based data, graph-structure, recommendation systems, structured reasoning etc. [24]. However, we can transform non-sequential data into sequential data through an offline preprocessing step [25]. and then use SSM for processing it.
Figure 1: (a) GPU (Nvidia A100) Latency of SSM models for recurrence vs convolutional methods when processing images with 4K sequence length and batch size of 64. (b) For convolution, the throughput per pixel decreases over sequence lengths, while the trend is flatter for the recurrence method.
@@
To meet these emerging computational demands, we propose EpochCore (ExPOnentially-Compressed History Core), a digital hardware accelerator designed to efficiently execute both structured SSM models (e.g., S4, Liquid-S4) and traditional dense neural networks (e.g., CNNs, RNNs, Transformers). EpochCore is built using an array of a novel Processing Element (PE) called LIMA-PE, which supports multiple MAC operation modes. These include recurrent integration with fixed (S4) and time-varying (Liquid-S4) coefficients, as well as standard GEMM dataflows for traditional DNN layers. LIMA-PE also has a unified design to process either real or complex number for the MAC operations. To reduce dynamic power, each LIMA-PE employs dual gated clocks to selectively minimize switching activity.
@@
    EpochCore Architecture: We introduce EpochCore, the first hardware accelerator that natively supports multiple Structured SSMs (e.g., S4, Liquid-S4), while retaining compatibility with traditional DNNs. This enables a unified inference system for both long-sequence and conventional models.
@@
We evaluate EpochCore on multiple LRA datasets, against two state-of-the-art accelerators: (1) a 1D systolic array optimized for long-kernel convolutions [3], and (2) a traditional SA with sparsity support [14], and GPUs. Against the 1D FFT-accelerated SA, EpochCore delivers on an average 25$\times$ speedup, 10$\times$ energy savings, and 30$\times$ lower memory bandwidth. For S4 and Liquid-S4 models on LRA datasets, EpochCore achieves up to 250$\times$ performance improvement and 45$\times$ energy reduction compared to traditional PE-based SAs. For inference latency on S4 layers, EpochCore outperforms GPUs by a factor of 2000$\times$. EpochCore also reduces memory bandwidth requirements by 30×30$\times$30 × compared to Boriakoff’s 1-D SA.
@@ II Background
In this section, we provide the background for the recently proposed Structured State-Space Sequential Models (S4). We also discuss the extension of S4 to include input-dependent time-varying coefficients, known as Liquid-S4. The mathematical framework for the computation of these models is formulated, highlighting the motivation for mapping them to hardware accelerators. These models rely on mapping inputs to internal state parameters, which are governed by a first-order ordinary differential equation (ODE).
@@ II-A State-Space Models (SSM)
Recent studies [35, 9, 10, 13] highlight the effectiveness of using the Scaled-Legendre (HiPPO-LegS) OP basis for improved accuracy. However, diagonalizing the coefficient matrix $\mathbf{A}$ for HiPPO-LegS can be computationally challenging. To address this, Gupta et al. [11] proposed the simpler Diagonal State Space (DSS) model, which computes the complex eigenvalues of the non-scaled HiPPO matrix. The eigenvalue-based diagonalization of the HiPPO matrix simplifies computation, making it an efficient and attractive approach.
@@
II-A1 Liquid S4
By introducing a dynamic, time-varying (or liquid) time-constant coefficient into the internal state dynamical equation (1), Hasani et al, demonstrated significant improvements in time-series prediction accuracy [13, 12]. This approach achieved SoTA results on LRA datasets. The updated dynamic equation for the internal state is expressed as:
@@ II-A1 Liquid S4
II-A2 Convolution Method for Solving SSMs
@@
Figure 2: (a) S4- [11], (b) Liq-S4-based [12] model with input sequences $u(t)$ and output sequences $y(t)$. Layer I determines the internal state space map recurrently. Layer II maps the internal state to the output sequence via a dense linear layer.
II-A3 Recurrent Method for Solving S4
@@
Hardware Accelerators: Mapping Layer I computations to GEMM operations on 2-D SAs, such as TPUs, is not straightforward. While the linear scaling $\mathbf{\overline{B}}\cdot u(t)$ can be implemented via diagonal matrix-vector multiplication, there are no native MAC operations for recurrent updates, as far as we know. To address this, we propose introducing a new cardinal MAC operation for recurrent updates, defined as:
@@
Layer II computations can be directly mapped to GEMM operations of a TPUs. The overall pipelining and throughput of the SSM model evaluation depend on the dataflow approach. In this paper, we recommend a variant of weight-stationary dataflow, which is particularly suitable for longer input and output sequences that move through the SA efficiently.
Figure 3: SSM-based models are multi-layered with interleaved SSM and DNN/CNN layers that to process input sequences $f(t)$ and generate output sequences or labels $g(t)$.
II-A4 Deep SSM
The SSM based models, illustrated in Figure 2, follows a defined sequence of operations outlined in Table II. Each SSM model can be configured to produce multi-dimensional output sequences, called heads ($H$), with each head trained to extract distinct features from the 1-D input sequence. These SSM model layers can be interleaved with DNN or CNN layers, as shown in Figure 3 to enhance their functionality.
Gu et al. [9] and Hasani et al. [13], have demonstrated SoTA accuracy for various LRA datasets by interleaving up to six SSM model layers. This multi-layer configuration, shown in Figure 3, enhances the model’s ability to process long-range dependencies. The EpochCore accelerator, discussed in section III, is designed to efficiently execute both SSM and non-SSM model layers such as DNN and CNNs, providing a versatile platform for diverse deep learning workloads.
@@ III EpochCore Architecture
Figure 4: (a) Architecture of the full System that uses EpochCore and (b) EpochCore micro-architecture. The EpochCore supports Epoch and GEMM operations.
@@ III-A EpochCore Micro-architecture
LRA datasets for Structured-SSM computations involve input and output sequences of length ranging from thousands to millions, making input-stationary (IS) or output-stationary (OS) approaches highly inefficient due to the need for extensive tiling. Since the weight matrix is the smallest among the weight, input or output matrices, a weight-stationary (WS) dataflow offers the lowest power consumption and highest throughput, while minimizing tiling needs. Hence, EpochCore adopts a modified WS dataflow for Structured-SSM computations. Preloading weights for an entire tile of size $n\times n$ requires significant on-chip SRAM bandwidth. EpochCore’s unified dataflow for evaluating Structured-SSM in a single SA computation allows weights to be stationary throughout the processing of a full batch of inputs, thereby amortizing the high on-chip SRAM bandwidth for loading weights. During Structured-SSM operations, the weights remain stationary while the 1D input sequence is broadcast across $n$ columns of the top row of the 2D SA, where $n$ is the size of the internal state-map of the S4 model.
@@
The LIMA-PEs in the SA are capable of performing MAC operations (for both S4s and traditional DNNs) and storing weights and the intermediate results from recurrent computations. The LIMA-PEs are interconnected, enabling data movement within the SA in West-to-East or North-to-South directions. Additionally, EpochCore supports diagonal data movement in the Northeast-to-Southwest direction, facilitating banded matrix multiplication and non-staggered multiplications required for Layer II (see Figure 2) computations in S4 models. The LIMA-PE is designed with a specific bit-precision. It supports data representation in either real or complex fixed-point formats. For complex values, the representation is achieved by multiplexing: the higher-order bits store the real part, while the lower-order bits store the imaginary part. As a result, processing complex values effectively halves the bit-precision compared to real-valued data.
@@
The controller unit decodes instructions from the host CPU to manage synchronized clocking, loading, and reset operations of the LIMA-PEs. To enable the novel programmable dataflow, ProDF, the controller configures specific modes for each LIMA-PE, facilitating the unique dataflow required to evaluate the S4 model for each tile. The three control input bits are multiplexed into the weight bus, with both the control bits and weights stored in the weight SRAM. The LIMA-PEs support both traditional MAC operations and specialized MAC operations needed for SSM recurrent computations. Each LIMA-PE contains two internal cache sets: one for stationary operands, such as the weight matrices and control inputs that remain static during SA operations, and another for intermediate computed values, such as state-map vectors, which may move dynamically through the SA. EpochCore execution involves the following sequential operating phases:
@@
Reset Phase: This phase initializes all the LIMA-PE by clearing all PE internal caches where we store the stationary and control inputs, and the intermediate ouput.
@@
Pre-Load Phase: During this phase, each LIMA-PE loads the control and weight data and is stored in the LIMA-PE stationary and control cache. These inputs, sourced from on-chip weight SRAM, prepare the PEs for subsequent operations.
@@
Compute Phase: In this phase, data flows through the PEs to evaluate the S4, Liquid-S4 or the traditional DNN model. The process is pipelined to enhance throughput. Depending on what is suitable for a model, one can use input stationary, weight stationary or output stationary dataflow in EpochCore.
@@
Readout Phase: In a traditional output stationary (OS) dataflow, the final output is stored in multiple rows of LIMA-PE. After the completion of the entire matrix computation, the outputs need to be sequentially read out onto the on-chip output SRAM during the readout phase.
@@ III-B LIMA-PE Micro-architecture
Figure 6: Data flow under various MAC operating modes of LIMA-PE. The green logic are part of traditional-PE design. LIMA-PE includes additional yellow logic for recurrent integration and banded matrix dataflows, and blue logic for time-varying integration support. The select inputs of the muxes, driven by the cached control inputs, enable the operating mode of the LIMA-PE. (a) and (b) are novel additions to the LIMA-PE to support recurrent integrations for S4 and Liquid-S4 models. (c) is a modified WS MAC operation to enable diagonal dataflow in the SA for banded-matrix multiplication and (d) support regular GEMM for OS data flow. WS & IS MAC operations, pass-through and sleep modes are not shown, but can be easily deduced.
The microarchitecture of the LIMA-PE is illustrated in Figure 5. Like traditional PEs, a LIMA-PE is single-buffered and incorporates registers to hold stationary and moving data. During the Pre-Load Phase, stationary data is loaded into the registers. No MAC operations are performed during the pre-load phase. In the subsequent Compute Phase, inputs and intermediate outputs are transferred across the SA as MAC operations process the data to compute final results.
For energy efficiency, the LIMA-PE generates two mutually exclusive internal clocks-Load Clock for Pre-Load phase and Compute Clock for Compute Phase, are generated in the clock controller circuit, ensuring the circuitry operates exclusively for its respective phase. While introducing clock-gating adds extra circuitry, it is a widely used technique for reducing power consumption. EpochCore includes additional control bits, which must be programmatically managed to enable the mutually exclusive clocks.
The throughput of an S4 layer processing in EpochCore is primarily limited by input and output bandwidth of the on-chip SRAM (more details in Section IV). Double buffering, a common method for improving performance, may not improve performance in EpochCore because loading weight matrices from weight SRAM to the LIMA-PEs typically accounts for only a small portion of total compute cycles. The sizes of the weight and input-output on-chip SRAM are 16MB each determined by the model size and these are chosen to minimize off-chip DRAM access.
@@
Accumulation Mode: In accumulation mode, the LIMA-PE performs MAC operations similar to traditional PEs, guided by cached input controls, as shown in Figure 5e. For the traditional GEMM operation in the WS, IS and OS dataflow, LIMA-PE allows data to flow from Result In to Result Out, i.e., North-to-South. For the recurrent and banded-matrix MAC operations, dataflow is enabled additionally from Data In to Data Out, i.e., Northeast-to-Southwest.

### v2

@@ Abstract
In this paper we introduce a specialized hardware accelerator,  EpochCore [^1] Epoch does not refer to training iterations, instead it is an acronym., for accelerating SSMs. EpochCore is based on systolic arrays (SAs) and is designed to enhance the energy efficiency and throughput of inference of SSM-based models for long-range sequence tasks. Within the SA, we propose a versatile processing element (PE) called LIMA-PE to perform traditional and specialized MAC operations to support traditional DNNs and SSMs. To complement the EpochCore microarchitecture, we propose a novel dataflow, ProDF, which enables highly efficient execution of SSM-based models. By leveraging the LIMA-PE microarchitecture and ProDF, EpochCore achieves on average 2000$\sim$ 2000$\times$ 2000 × improvement in performance on LRA datasets compared to a GPU and 250×~250$\times$250 × gains in performance and 45×~45$\times$45 × improvement in energy efficiency, over traditional SA-based accelerators (TPU).
@@ I Introduction
To address this limitation, recent research has explored alternative architectures based on State Space Models (SSMs) [25], which offer a promising framework for scalable sequence modeling. The Linear State Space Layer (LSSL) [10] demonstrated early improvements over both RNNs and Transformers on simpler sequence tasks. Further enhancements to SSMs introduced structured state transition matrices—e.g., in S4 [9], S5 [27], DSS [11], and H3 [6]—to boost both accuracy and efficiency. More recent models, such as Liquid-S4 [13, 12], GSS [23], and Mamba [7, 21], integrate time-varying parameters and gating mechanisms, achieving state-of-the-art results on long-range benchmarks such as the Long Range Arena (LRA) [29] as shown in Table I. These benchmarks encompass a wide range of applications, including 500+ token language tasks, 1000+ time steps for time-series data, and hundreds of video frames. For example, Gu et al. [9] report that S4 achieves a 5.19$\times$ speedup and 0.091$\times$ memory usage compared to Transformers on 4K-length inputs, while also delivering over 60$\times$ throughput gains in token processing. SSM is not suited for non-sequential data such as grid-based data, graph-structure, recommendation systems, structured reasoning etc. [24]. However, we can transform non-sequential data into sequential data through an offline preprocessing step [25]. and then use SSM for processing it.
Figure 1: (Left) GPU (Nvidia A100) latency of SSM models for recurrence vs convolutional methods when processing images with 4K sequence length and batch size of 64. (Right) For convolution, the throughput per pixel decreases over sequence lengths.
@@
To meet these emerging computational demands, we propose EpochCore (ExPOnentially-Compressed History Core), a digital hardware accelerator designed to efficiently execute both structured SSM models (e.g., S4, Liquid-S4) and traditional dense neural networks (e.g., CNNs, RNNs, Transformers). EpochCore is built using an array of a novel Processing Element (PE) called LIMA-PE, which supports multiple MAC operation modes. These include recurrent integration with fixed (S4) and time-varying (Liquid-S4) coefficients, as well as standard GEMM dataflows for traditional DNN layers. LIMA-PE also has a unified design to process either real or complex numbers for the MAC operations. To reduce dynamic power, each LIMA-PE employs dual gated clocks to selectively minimize switching activity.
@@
    EpochCore Architecture: We introduce EpochCore, the first hardware accelerator that natively executes multiple Structured SSMs (e.g., S4, Liquid-S4), while retaining support for traditional DNNs. Effectively, we have a unified inference system for both long-sequence and conventional models.
@@
We evaluated EpochCore on multiple LRA datasets, comparing it against three SoTA general-purpose accelerators: (1) a 1D systolic array optimized for long-kernel convolutions [3], (2) a traditional SA with sparsity support [14], and (3) GPUs. Compared to the 1D FFT-accelerated SA, run on S4 and Liquid-S4 models, EpochCore delivers on an average 25$\times$ speedup, 10$\times$ energy savings, and 30$\times$ lower memory bandwidth. For S4 and Liquid-S4 models on LRA datasets, EpochCore achieves up to 250$\times$ performance improvement and 45$\times$ energy reduction compared to traditional PE-based SAs. For inference latency on S4 layer and Liquid-S4 layers, EpochCore outperforms GPUs by a factor of 2000$\times$.
We also compared EpochCore’s performance improvements over GPUs against three recently proposed SSM accelerators: (1) For H3 [6] model inference, EpochCore achieves 3860×3860$\times$3860 × improvement, while the VGA [17] accelerator shows a 4×4$\times$4 × improvement over a GPU. (2) For Mamba model inference, EpochCore provides a 4.75×4.75$\times$4.75 × improvement, while on Marca [19] achieves an 11.66×11.$6$ and FastMamba [34] achieves a 6.06×6.06$\times$6.06 × speedup over a GPU.
@@ II Background
In this section, we provide the background for the recently proposed Structured State-Space Sequential Models (S4). We also discuss the extension of S4 to include input-dependent time-varying coefficients, known as Liquid-S4. These models rely on mapping inputs to internal state parameters, which are governed by a first-order ordinary differential equation (ODE).
@@ II-A State-Space Models (SSM)
Recent studies [35, 9, 10, 13] highlight the effectiveness of using the Scaled-Legendre (HiPPO-LegS) OP basis for improved accuracy. However, diagonalizing the coefficient matrix $\mathbf{A}$ for HiPPO-LegS can be computationally challenging. To address this, Gupta et al. [11] proposed the simpler Diagonal State Space (DSS) model, which computes the complex eigenvalues of the non-scaled HiPPO matrix. The eigenvalue-based diagonalization of the HiPPO matrix simplifies computation, making it an efficient and attractive approach as shown by Gu et al in S4  [9].
@@
Liquid-S4: By introducing a dynamic, time-varying (or liquid) time-constant coefficient into the internal state dynamical equation (1), Hasani et al, demonstrated significant improvements in time-series prediction accuracy [13, 12]. The updated dynamic equation for the internal state is expressed as:
@@
II-A1 Convolution Method for Solving SSMs
@@
Figure 2: (a) S4- [11], (b) Liq-S4-based [12] model with input sequences $u(t)$ and output sequences $y(t,H)$. Layer I determines the internal state space map recurrently. Layer II maps the internal state to the output sequence via a dense linear layer.
II-A2 Recurrent Method for Solving S4
@@
Hardware Accelerators: 1 Layer I computations to GEMM operations on 2-D SAs, such as TPUs, is not straightforward. While the linear scaling $\mathbf{\overline{B}}\cdot u(t)$ can be implemented via diagonal matrix-vector multiplication, there are no native MAC operations for recurrent updates. To address this, we introduce a new cardinal MAC operation for recurrent updates, defined as:
@@
Layer II computations can be directly mapped to GEMM operations of a TPU. The overall pipelining and throughput of the SSM model evaluation depend on the dataflow approach. In this paper, we recommend a variant of weight-stationary dataflow, which is particularly suitable for longer input and output sequences that move through the SA efficiently.
Figure 3: SSM-based models are multi-layered with interleaved SSM and DNN/CNN layers that process input sequences $f(t)$ and generate output sequences or labels $g(t)$.
Deep SSM: The SSM based models, illustrated in Figure 2, follows a defined sequence of operations outlined in Table II. Each SSM model can be configured to produce multi-dimensional output sequences, called heads ($H$), with each head trained to extract distinct features from the 1-D input sequence. These SSM model layers can be interleaved with DNN or CNN layers, as shown in Figure 3 to enhance their functionality.
Gu et al. [9] and Hasani et al. [13], have demonstrated SoTA accuracy for various LRA datasets by interleaving up to six SSM model layers with DNN layers. This multi-layer configuration, shown in Figure 3, enhances the model’s ability to process long-range dependencies. The EpochCore accelerator, discussed in section III, is designed to efficiently execute both SSM and non-SSM model layers such as DNN and CNNs, providing a versatile platform for diverse deep learning workloads.
@@ III EpochCore Architecture
Figure 4: (a) Architecture of the full System that uses EpochCore and (b) EpochCore micro-architecture. The EpochCore supports SSM and GEMM operations.
@@ III-A EpochCore Micro-architecture
LRA datasets for Structured-SSM (S4 and Liquid S4) computations involve input and output sequences of length ranging from thousands to millions, making input-stationary (IS) or output-stationary (OS) approaches highly inefficient due to the need for extensive tiling. Since the weight matrix is the smallest among the weight, input or output matrices, a weight-stationary (WS) dataflow offers the lowest power consumption and highest throughput, while minimizing tiling needs. Hence, EpochCore adopts a modified WS dataflow for Structured-SSM computations. The modification is in programatically altering the flow of output through the processing elements while keeping the weights stationary. Preloading weights for an entire tile of size $n\times n$ requires significant on-chip SRAM bandwidth. EpochCore’s unified dataflow for evaluating Structured-SSM in a single SA computation allows weights to be stationary throughout the processing of a full batch of inputs, thereby amortizing the high on-chip SRAM bandwidth for loading weights. During Structured-SSM operations, the weights remain stationary while the 1D input sequence is broadcast across $n$ columns of the top row of the 2D SA, where $n$ is the size of the internal state-map of the S4 model.
@@
The LIMA-PEs in the SA are capable of performing MAC operations (for both S4s and traditional DNNs) and storing weights and the intermediate results from recurrent computations. The LIMA-PEs are interconnected, enabling data movement within the SA in West-to-East or North-to-South directions. Additionally, EpochCore supports diagonal data movement in the Northeast-to-Southwest direction, facilitating banded matrix multiplication and non-staggered multiplications required for Layer II (see Figure 2) computations in S4 models. The LIMA-PE is designed with a specific bit-precision. It supports data representation in either real or complex fixed-point formats. For complex values, the representation is achieved by sharing the higher-order bits to store the real part, while using lower-order bits store the imaginary part. As a result, processing complex values effectively halves the bit-precision compared to real-valued data.
@@
The controller unit decodes instructions from the host CPU to manage synchronized clocking, loading, and reset operations of the LIMA-PEs. To enable the novel programmable dataflow, ProDF (discussed later in this section), the controller configures specific modes for each LIMA-PE, facilitating the unique dataflow required to evaluate the S4 model for each tile. The three control input bits are multiplexed into the weight bus, with both the control bits and weights stored in the weight SRAM. The LIMA-PEs support both traditional MAC operations and specialized MAC operations needed for SSM recurrent computations. Each LIMA-PE contains two internal buffers sets, implemented as registers: one for stationary operands, such as the weight matrices and control inputs that remain static during SA operations, and another for intermediate computed values, such as state-map vectors, which may move dynamically through the SA. EpochCore execution involves the following sequential operating phases:
@@
Reset Phase: This phase initializes all the LIMA-PE by clearing all PE internal buffers, where we store the stationary and control inputs, and the intermediate outputs.
@@
Pre-Load Phase: During this phase, each LIMA-PE loads the control and weight data into the LIMA-PE control and stationary buffers, respectively. These inputs, sourced from on-chip weight SRAM, prepare the PEs for subsequent operations.
@@
Compute Phase: In this phase, data flows through the PEs to evaluate the S4, Liquid-S4 or the traditional DNN model. The process is pipelined to enhance throughput. Depending on what is suitable for the S4, Liquid-S4 or DNN model, one can use input stationary, weight stationary or output stationary dataflow in EpochCore.
@@
Readout Phase: The final output is stored in multiple rows of LIMA-PE. After the completion of the entire matrix computation, the outputs need to be sequentially read out onto the on-chip output SRAM during the readout phase.
@@ III-B LIMA-PE Micro-architecture
Figure 6: Data flow under various MAC operating modes of LIMA-PE. The green logic are part of traditional-PE design. LIMA-PE includes additional yellow logic for recurrent integration and banded matrix dataflows, and blue logic for time-varying integration support. The select inputs of the muxes, driven by the buffered control inputs, enable the operating mode of the LIMA-PE. (a) and (b) are novel additions to the LIMA-PE to support recurrent integrations for S4 and Liquid-S4 models. (c) is a modified WS MAC operation to enable diagonal dataflow in the SA for banded-matrix multiplication and (d) support regular GEMM for OS data flow. WS & IS MAC operations, pass-through and sleep modes are not shown, but can be easily deduced.
The microarchitecture of the LIMA-PE is illustrated in Figure 5. Like traditional PEs, a LIMA-PE is single-buffered and incorporates registers to hold stationary data in the stationary buffer and moving data in the output buffer of the compute unit. During the Pre-Load Phase, stationary data is loaded into the registers. No MAC operations are performed during the pre-load phase. In the subsequent Compute Phase, inputs and intermediate outputs are transferred across the SA as MAC operations process the data to compute final results.
For energy efficiency, the LIMA-PE generates two mutually exclusive internal clocks-Load Clock for Pre-Load Phase and Compute Clock for Compute Phase, in the clock controller circuit. While introducing clock-gating adds extra circuitry, it is a widely used technique for reducing power consumption. EpochCore includes additional control bits, which must be programmatically managed to enable the mutually exclusive clocks.
The throughput of an S4 layer processing in EpochCore is primarily limited by input and output bandwidth of the on-chip SRAM (more details in Section IV). Double buffering, a common method for improving performance, may not improve performance in EpochCore because loading weight matrices from weight SRAM to the LIMA-PEs typically accounts for only a small portion of total compute cycles. The sizes of the weight and input-output on-chip SRAM are 16MB, each determined by the S4 model size, and these are chosen to minimize off-chip DRAM access.
@@
Accumulation Mode: In accumulation mode, the LIMA-PE performs MAC operations similar to traditional PEs, guided by buffered input controls, as shown in Figure 5e. For the traditional GEMM operation in the WS, IS and OS dataflow, LIMA-PE allows data to flow from Result In to Result Out, i.e., North-to-South. For the recurrent and banded-matrix MAC operations, dataflow is enabled additionally from Data In to Data Out, i.e., Northeast-to-Southwest.